OpenStudy (askme12345):

find the inverse of f^-1(x)

OpenStudy (askme12345):

f(x)=2x-4/5x+6

OpenStudy (abb0t):

solve for \(\sf \color{red}{x}\) in terms of \(\sf \color{orange}{y}\)

OpenStudy (askme12345):

so would that be x=2y-4/5y+6

OpenStudy (abb0t):

Yessir.

OpenStudy (askme12345):

is that the final answer? would you be able to assist me with one more, f(x)=3−ln5x+20

OpenStudy (abb0t):

What?! Where did you get natural log from?!?!?!

OpenStudy (abb0t):

And no, that is not the asnwer. You need to SOLVE for \(\sf \color{red}{x}\)!

OpenStudy (askme12345):

that is a second problem i need assistance with

OpenStudy (askme12345):

i am not sure how to complete solving for x for the first one

OpenStudy (abb0t):

Well, get rid of the demonimator first. Multiply both sides by \(\sf \color{red}{5y+6}\)

OpenStudy (askme12345):

okay so your solution is x=10y^2-8y-24

OpenStudy (abb0t):

You get: x(5x+6)=2y-4 Now, distribute the \(\sf \color{blue}{x}\) across the parenthesis. 5xy+6x=2y-4 Next, group like terms together. So subtract 2y from both sides: 5xy-2y+6=-4 subtract by 6 to both sides: 5xy-2y = -10 factor out the \(\sf \color{orange}{y}\): y(5x-2)=-10 and divide to isolate y! the end.

OpenStudy (abb0t):

Does that make sense?

OpenStudy (askme12345):

Yes! So Y=-10/5x-2

OpenStudy (askme12345):

Thank you so much!! Superb answer

OpenStudy (abb0t):

Correct. Best of luck!

OpenStudy (askme12345):

Thank you. Would you be able to assist me with finding the inverse of f(x)=\[\sqrt{3x+5}\] I understand i must substitute all of the x's for y's so...x= \[\sqrt{3y+5}\] so x^2 = 3y+5? Is that where i would move on too? Im not sure where to go from there